OpenStudy (anonymous):

Calculus: determine the power series expansion for 1/(x^4 + 4) and its radius of convergence?

OpenStudy (anonymous):

Well you could use something similar to the geometric series form: \[\sum_{n=0}^{\infty}ar^n=1/(1-r) \iff |r|<1. \] You can rewrite this as such I believe.

OpenStudy (anonymous):

That should read =a/(1-r) sorry about that.

OpenStudy (anonymous):

right right. i get ya. rearranging it to be something like \[1/4(1-(-1/4x^4))\]? where the radius of convergence would be \[\left| -1/4x^4 \right| < 1\] there for \[\left| x \right|<\sqrt{2}\]

OpenStudy (anonymous):

Thats what I got :D

OpenStudy (anonymous):

Sweet thanks :)